# Super LiDAR Reflectance for Robotic Perception

### v1

@@ III-B Static-to-Dynamic Domain Gap
The above ground-truth supervision data collection is collected when the LiDAR is stationary. However, the proposed reflectance densification method is generally intended for robot perception tasks when the LiDAR is in-motion. In practical motion mode, it is necessary to fuse via odometry estimation a certain number of consecutive scans to construct a relatively dense reflectance image as network input. When fusing, using too few scans (e.g., only one or two scans) often results in excessively sparse data, leading to lacking sufficient information. On the other hand, fusing too many scans can introduce errors caused by odometry estimation inaccuracies or noise points generated by dynamic objects in the scene. To address the influence of dynamic objects in such scenarios, we adopt the preprocessing method proposed in [31] to filter the raw point cloud data, effectively removes motion artifacts caused by dynamic objects and mitigating their adverse effects on the fusion process. In our case, the network input is generated by fusing $N$ (a small number) consecutive scans using odometry estimates obtained by [15]. By comparing with the static ground-truth training data generation, there obviously exists a gap between the static training data and dynamic inference ones, which can degrade performance during motion scenario.
@@
Figure 5: Comparison of reflectance images fused from 5 scans in stationary and in-motion cases. The image shows the corresponding reflectance images. In the stationary case, the 5-scan inputs are aligned, whereas the in-motion case reveals irregularities and misalignment due to motions.
Mathematically, the input in the static training setup can be represented as
@@
where $\mathcal{P}()$ represents the projection operation mapping the 3D point cloud to a 2D reflectance image, $\mathbf{R}_{i}(x,y,z)$ is the raw point cloud of the $i$-th scan, $\sum$ denotes the geometric merging of multi-scan point clouds. In contrast, the input to the training stage (also for the inference stage) when the robot is in motion is represented as
@@ III-C Adaptive Calibration and Densification Network
Figure 6: Framework of the proposed method for LiDAR reflectance densification. The framework includes an Adaptive Feature Module (AFM) for feature extraction, an Encoder-Decoder structure for spatial reconstruction, and a Dynamic Compensation Module (DCM) to correct intensity.
@@
Considering that convolutional neural networks (CNNs) effectively extract spatial features with computational complexity approximately proportional to the number of input elements $\mathcal{O}(n)$ and maintain efficiency for large images, they were chosen as the backbone of our densification network. In contrast, Transformers[32], despite their strong modeling capabilities, exhibit a quadratic complexity of $\mathcal{O}(n^{2})$ with respect to the input size, which significantly limits their application in resource-constrained scenarios, especially for high-resolution images[33, 34]. Similarly, diffusion models[35], while demonstrating impressive performance in image restoration, rely on iterative processes to generate the final output, resulting in long inference times[36], making them hard for real-time robotics applications. Fig. 6 demonstrates the structure of our densification network, which employs a U-shaped architecture[37] where an Adaptive Fusion Module (AFM) and a Dynamic Compensation Module (DCM) are proposed to achieve densification. These modules address two key challenges: 1) adaptively fusing multi-scale information from non-uniform and sparse LiDAR scanning points and 2) mitigating the intensity decay that occurs with distance and incidence angle.
@@
Let $L_{1}\in\mathbb{R}^{H\times W}$ and $D_{1}\in\mathbb{R}^{H\times W}$ are the LiDAR reflectance and depth images as the network input. The depth information is introduced in the reflectance image reconstruction during the reflectance compensation step. To maximize the utilization of cross-modal interaction and reduce model redundancy, $L_{1}$ and $D_{1}$ are first concatenated along the channel dimension as input to the network. Then, these inputs are processed through an encoder-decoder structure integrated with AFM to produce preliminary reconstructions of dense reflectance images $L_{2}\in\mathbb{R}^{H\times W}$ and depth images $D_{2}\in\mathbb{R}^{H\times W}$, where the first and second channels of the decoder are $L_{2}$ and $D_{2}$, respectively. Furthermore, to eliminate artifacts caused by signal intensity decay, the DCM is applied to compensate for these effects, resulting in the final dense LiDAR reflectance image $L{3}\in\mathbb{R}^{H\times W}$. The DCM employs a deep learning-based dynamic compensation model to avoid the complexity of manually designing the parameters required in traditional methods while maintaining end-to-end reconstruction.
@@
Encoder and Decoder. The encoder has four stacked blocks, each comprising two sequentially connected layers with ReLU activation followed by an average pooling layer for downsampling. The convolutional kernels are set to a size of $3\times 3$. The pooling window and stride of the pooling layer are configured with 2 for downsampling by a factor of 2. The decoder has a symmetric structure, comprising four decoder blocks. Within each decoder block, the input feature map is upsampled using a transposed convolution module. Subsequently, two $3\times 3$ convolutions are applied for local information fusion. To enhance multiscale feature integration, we implement two strategies: 1) an AFM module is inserted between the encoder and decoder to facilitate multi-scale spatial information; 2) similar to[37], features from the encoder are incorporated into the decoder. Corresponding feature maps of identical spatial dimensions from four blocks are concatenated to integrate multi-level information.
@@
Adaptive Fusion Module. Due to the scanning characteristics of NRS-LiDAR, reflectance images captured in short durations are typically sparse and unevenly distributed. Besides, since objects in scenes possess multiscale attributes and vary in distance, fusing raw LiDAR reflectance data for incomplete information densification demands that the model possess adaptive capabilities and the ability to fuse multiscale structural features. Traditional convolutional kernels with limited receptive fields struggle to capture information across large regions, and larger kernels significantly increase computational cost. Inspired by[38, 39], we utilize parallel dilated convolutions and deformable convolutions to adaptively integrate multiscale information (Fig. 6). Specifically, we employ multiple dilation rates in parallel dilated convolutions to expand the receptive field with minimal additional computational cost, thereby enhancing the model’s capacity for multiscale feature extraction. Simultaneously, we incorporate a deformable convolution to facilitate adaptive information fusion based on local contexts, thereby augmenting the model’s ability to extract detailed information. The outputs from these parallel convolutions are subsequently concatenated along the channel dimension to form the output features of the AFM module.
@@
Figure 7: Comparison between the raw dense reflectance and the compensated result. The compensated result maintains more consistent reflectance values for objects with the same surface material (yellow boxes in (a) and (b)) and exhibits clearer texture details (blue boxes in (c) and (d)).
@@
Dynamic Compensation Module. The LiDAR intensity is primarily influenced by three factors: the distance traveled by light, the object’s surface reflectance, and the incident angle of light. In general, it is assumed that the reflectance of the object’s surface remains relatively consistent. Therefore, compensating for the effects of distance and incident angle is critical for intensity measurement[30]. As introduced in[30], the captured intensity can be expressed as
@@
where $I_{e}$ represents the laser’s emission power, $R$ denotes the distance from the LiDAR to the object’s surface, $\alpha$ is the incident angle, and $\rho$ indicates the surface reflectance of the object. The term $\eta(R)$, which accounts for the near-distance effect, can be expressed as
@@
where $C$ is a proportionality constant that accounts for system-specific calibration factors. $I_{c}$ represents the corrected or true intensity.
Figure 8: Qualitative comparison of reflectance image densification across different methods in indoor and outdoor scenarios. IP-basic[40], which relies on non-learning techniques, struggles to recover structures in regions with large holes, resulting in distorted outputs. Furthermore, Sparse-to-Dense[41] and pNCNN[42] exhibit noticeable deficiencies in both smoothness and edge preservation. Our method with AFM demonstrates superior performance in preserving edge structures and enhancing details. The compensation results show that the intensity values for the same surface material (e.g., the same wall) are maintained more consistently (as shown in yellow boxes).
@@
In practice, we collected data for objects with the same material under controlled angles and distances, fitted the required parameters by Eq. 8, and applied them to the raw dense LiDAR reflectance images to generate supervised compensated results. The compensated results are shown in Fig. 7. However, the above methods are further limited by data collection constraints and sensor configurations, making them inadequate for capturing complex, non-linear relationships in diverse environments. To address this, we use a simplified, parameterized compensation model, as shown in Eq. (9) and (10), which is based on Eq. (8) but incorporates an additional decay factor to comprehensively handle both near-distance and far-distance effects. For distance compensation, the distance correction factor $\frac{R^{2}}{\eta(R)}$ becomes significantly large when $R$ is large. Therefore, we approximated it using a second-order polynomial combined with an exponential decay factor to handle far-distance effects.
$g(R)=\frac{\left(\alpha R^{2}+\beta R+\gamma\right)\cdot\exp(-\lambda R)}{\eta(R)},$ (9)
where $\alpha$, $\beta$, $\gamma$, and $\lambda$ are learnable parameters predicted dynamically by the network to capture the non-linear relationship between intensity and range.
For incidence-angle compensation, the intensity is normalized by $\cos(\alpha)$, following Lambert’s cosine law[44]. By combining both distance and angle compensation, the corrected intensity $I_{\text{com}}$ is expressed as
$I_{\text{com}}=\frac{I_{\text{m}}}{g(R)\cdot\cos(\alpha)\cdot\kappa},$ (10)
where $I_{\text{m}}$ represents the measured intensity, equal to the $I(R,\alpha,\rho)$ in Eq.6, and $\kappa$ is a learnable parameter to account for variations in surface reflectance and sensor-specific factors. In our method, convolutional layers are used to dynamically predict the compensation parameters $\alpha$, $\beta$, $\gamma$, $\lambda$ and $\kappa$ for each pixel based on local context, the distance $R$ and the incident angle $\alpha$ are both provided by $D_{2}$. These parameters are then applied to $I_{\text{m}}$ using Eq. (10), enabling the model to perform intensity compensation.
IV Experiments
We conducted a series of experiments to evaluate the effectiveness and robustness of the proposed method, focusing on the real-time LiDAR reflectance image densification and its applications in loop closure detection and traffic lane detection. Data collection and real-world experiments were conducted using a Livox MID-360, mounted on a Giraffe ground robot equipped with an NVIDIA RTX 3090 GPU and an Intel i7-1165G7 CPU. The experiments covered diverse scenarios, including varying lighting conditions (daylight, overcast, nighttime) and urban environments, to comprehensively evaluate the proposed method.
IV-A Evaluation of LiDAR Reflectance Densification
@@
Since there are limited public datasets or few prior works specifically addressing reflectance densification, we trained several depth completion models and included traditional interpolation methods for comparison[41, 42, 40, 43]. Additionally, an ablation study was conducted to analyze the contribution of the Adaptive Fusion Module (AFM). Table I presents the quantitative results, while Fig. 8 demonstrates the qualitative superiority of our method. Our method achieves high accuracy (highest PSNR/SSIM, lowest RMSE/MAE) while maintaining competitive runtime. Notably, as Marigold-dc[43] is a zero-shot diffusion-based method designed for depth completion, its performance is less effective in our task context; therefore, we only provide its quantitative results.
To evaluate the performance of our method in generating dense reflectance images during robot motion, we tested on data (5-frame accumulation) captured when robot was subject to aggressive motion poses on uneven road. The resulting PSNR values (25.75–26.82) demonstrate the robustness of our approach against alignment noise and non-linear motion artifacts (Fig. 9). Additionally, we compared real-time results at various robot speeds in the supplementary material.
@@ IV-B1 Loop Closure Detection
To validate the effectiveness of our method, we adopt the approach of[2] but replace the high-resolution reflectance data acquired by high-cost LiDAR with the dense reflectance images generated by our method using Livox MID-360. Specifically, [2] extracts ORB feature descriptors[47] from reflectance images for loop closure detection. These descriptors are then represented as bag-of-words vectors for querying loop closures. After retrieving a candidate match, ORB descriptor matching is performed, and outliers are removed using PnP RANSAC[48] for verification. We conducted experiments in three distinct environments under two lighting conditions: daytime and nighttime (Fig. 10), using data collected under varying conditions to assess the generalization capability of our approach. Because the adopted loop closure detection method[2] is not based on deep learning, for a fair comparison, we compared our loop closure detection results with other non-deep-learning methods, including [18, 46], which are representative LiDAR-descriptor based methods. The results in Table II demonstrate that our method consistently outperforms competing approaches, maintaining reliable performance despite environmental and illumination changes.

### v2

@@ III-B Static-to-Dynamic Domain Gap
The above ground-truth supervision data collection is collected when the LiDAR is stationary. However, the proposed densification method is generally intended for robot perception tasks when the LiDAR is in-motion. In practical motion mode, it is necessary to fuse via odometry estimation a certain number of consecutive scans to construct a relatively dense intensity image as network input. When fusing, using too few scans (e.g., only one or two scans) often results in excessively sparse data, leading to lacking sufficient information. On the other hand, fusing too many scans can introduce errors caused by odometry estimation inaccuracies or noise points generated by dynamic objects in the scene. To address the influence of dynamic objects in such scenarios, we adopt the preprocessing method proposed in [47] to filter the raw point cloud data, effectively removes motion artifacts caused by dynamic objects and mitigating their adverse effects on the fusion process. In our case, the network input is generated by fusing $N$ consecutive scans using odometry estimates obtained by [50]. We empirically selected $N=5$ to balance input density against motion-induced artifacts (see Supplementary Material). By comparing with the static ground-truth training data generation, there obviously exists a gap between the static training data and the dynamic inference ones, which can degrade performance during motion scenarios.
@@
Figure 5: Comparison of intensity images fused from 5 scans in stationary and in-motion cases. The image shows the corresponding intensity images. In the stationary case, the 5-scan inputs are aligned, whereas the in-motion case reveals irregularities and misalignment due to motions.
The input in the static training setup can be represented as:
@@
where $\mathcal{P}(\cdot)$ represents the projection operation from 3D point cloud to 2D intensity image, $\mathbf{R}_{i}(x,y,z)$ is the raw point cloud of the $i$-th scan, $\sum$ denotes the geometric merging of multi-scan point clouds. In contrast, the input to the training stage (also the inference stage) when the robot is in motion is:
@@ III-C Adaptive Calibration and Densification Network
Figure 6: Framework of the proposed method for LiDAR intensity image densification. The framework includes an Adaptive Feature Module (AFM) for feature extraction, an Encoder-Decoder structure for spatial reconstruction, and a Dynamic Compensation Module (DCM) to correct intensity.
@@
Considering that convolutional neural networks (CNNs) effectively extract spatial features with computational complexity approximately proportional to $\mathcal{O}(n)$ while maintaining efficiency, they were chosen as the backbone of our densification network. In contrast, Transformers [39] exhibit quadratic complexity $\mathcal{O}(n^{2})$, limiting their application in resource-constrained scenarios [52, 22]. Similarly, diffusion models [15] rely on iterative processes, leading to long inference times [21] that hinder real-time applications. Fig. 6 shows our densification network, which employs a U-shaped architecture [33] with an Adaptive Fusion Module (AFM) and a Dynamic Compensation Module (DCM) to achieve densification. These modules address two key challenges: 1) adaptively fusing multi-scale information from sparse LiDAR points and 2) mitigating intensity decay with distance and incidence angle.
@@
Let $L_{1}\in\mathbb{R}^{H\times W}$ and $D_{1}\in\mathbb{R}^{H\times W}$ be the LiDAR intensity and depth images as input into the network. Depth information is introduced in the intensity image reconstruction during the intensity compensation step. To maximize the utilization of cross-modal interaction and reduce model redundancy, $L_{1}$ and $D_{1}$ are first concatenated along the channel dimension as input to the network. Then, these inputs are processed through an encoder-decoder structure integrated with AFM to produce preliminary reconstructions of dense intensity images $L_{2}\in\mathbb{R}^{H\times W}$ and depth images $D_{2}\in\mathbb{R}^{H\times W}$, where the first and second channels of the decoder are $L_{2}$ and $D_{2}$, respectively. Furthermore, to eliminate artifacts from the signal intensity decay, DCM is applied to compensate for these effects, resulting in the final dense intensity image $L{3}\in\mathbb{R}^{H\times W}$. DCM employs a learning-based dynamic compensation model to avoid manually designing the parameters required in traditional methods while maintaining end-to-end reconstruction.
@@
Encoder and Decoder. The encoder comprises four stacked blocks. Each block contains two sequentially connected $3\times 3$ convolutional layers with ReLU activation, followed by an average pooling layer that downsamples by a factor of 2. Symmetrically, the decoder consists of four blocks. In each decoder block, the input feature map is first upsampled via a transposed convolution module, and then processed by two $3\times 3$ convolutions for local information fusion. To enhance multi-scale feature integration, we adopt two strategies: 1) inserting an AFM module between the encoder and decoder to facilitate multi-scale spatial information; and 2) following the U-Net architecture [33], concatenating corresponding feature maps of identical spatial dimensions from the four encoder blocks into the decoder.
@@
Adaptive Fusion Module. Due to the scanning characteristics of NRS-LiDAR, intensity images captured in short durations are typically sparse and unevenly distributed. Besides, since objects in scenes possess multiscale attributes and vary in distance, fusing raw LiDAR intensity data for incomplete information densification demands that the model possess adaptive capabilities and the ability to fuse multiscale structural features. Traditional convolutional kernels with limited receptive fields struggle to capture information across large regions, and larger kernels significantly increase computational cost. Inspired by[2, 17], we utilize parallel dilated convolutions and deformable convolutions to adaptively integrate multiscale information (Fig. 6). Specifically, parallel dilated convolutions with varying dilation rates expand the receptive field with minimal computational overhead, enhancing multi-scale feature extraction. Simultaneously, deformable convolutions adaptively fuse information based on local contexts, augmenting the model’s ability to extract fine details. Outputs from these parallel convolutions are concatenated along the channel dimension to form the AFM module’s output features.
@@
Figure 7: Comparison between the raw dense intensity and the compensated result. The compensated result maintains more consistent intensity values for objects with the same surface material (yellow boxes in (a) and (b)) and exhibits clearer texture details (blue boxes in (c) and (d)).
@@
Dynamic Compensation Module. The LiDAR intensity is primarily influenced by three factors: the distance traveled by light, the object’s surface reflectance, and the incident angle of light. In general, it is assumed that the reflectance of the object’s surface remains relatively consistent. Therefore, compensating for the effects of distance and incident angle is critical for intensity measurement[8]. As introduced in[8], the captured intensity can be expressed as:
@@
where $I_{e}$ represents the laser’s emission power, $R$ denotes the distance from the LiDAR to the object’s surface, $\alpha$ is the incident angle, and $\rho$ indicates the surface reflectance of the object. The term $\eta(R)$, which accounts for the near-distance effect, can be expressed as:
@@
where $C$ is a proportionality constant that accounts for system-specific calibration factors. $I_{c}$ is the corrected or true intensity.
Figure 8: Qualitative comparison of intensity image densification across different methods in indoor and outdoor scenarios. IP-basic [18], a non-learning method, struggles to recover structures in regions with large holes, resulting in distorted outputs. Sparse-to-Dense [26] and pNCNN [7] exhibit noticeable deficiencies in both smoothness and edge preservation. Our method with AFM demonstrates superior performance in preserving edge structures and enhancing details. The compensation results show that the intensity values for the same surface material (e.g., the same wall) are maintained more consistently (yellow boxes).
In practice, we collected data for objects with the same material under controlled angles and distances, fitted the required parameters by Eq. 8, and applied them to generate supervised compensated results for training. The compensated results are shown in Fig. 7. However, the above methods are further limited by data collection constraints and sensor configurations, making them inadequate for capturing complex, non-linear relationships in diverse environments. To address this, we use a simplified, parameterized compensation model, as shown in Eq. (9) and (10), which is based on Eq. (8) but incorporates an additional decay factor to comprehensively handle both near-distance and far-distance effects. For distance compensation, the distance correction factor $\frac{R^{2}}{\eta(R)}$ becomes significantly large when $R$ is large. Therefore, we approximated it using a second-order polynomial combined with an exponential decay factor to handle far-distance effects.
$g(R)=\frac{\left(\xi R^{2}+\beta R+\gamma\right)\cdot\exp(-\lambda R)}{\eta(R)},$ (9)
where $\xi$, $\beta$, $\gamma$, and $\lambda$ are learnable parameters predicted dynamically by the network to capture the non-linear relationship between intensity and range.
For incidence-angle compensation, the intensity is normalized by $\cos(\alpha)$, following Lambert’s cosine law[16]. By combining both distance and angle compensation, the corrected intensity $I_{\text{com}}$ is expressed as:
$I_{com}=\frac{I_{m}\cdot g(R)}{\cos(\alpha)\cdot\kappa}$ (10)
where $I_{\text{m}}$ represents the measured intensity, equal to the $I(R,\alpha,\rho)$ in Eq.6, and $\kappa$ is a learnable parameter to account for deviations from ideal Lambertian reflectance and sensor-specific factors. In our method, we employ a lightweight CNN comprising three convolutional layers with $3\times 3$ kernels to dynamically predict per-pixel compensation parameters ($\xi$, $\beta$, $\gamma$, $\lambda$, $\kappa$) from local context. The propagation distance $R$ is obtained from the densification module’s output depth map $D_{2}$. To compute the incidence angle $\alpha$, we first back-project $D_{2}$ into 3D to recover surface points, estimate per-pixel surface normals via local PCA plane fitting, and then use the LiDAR beam direction $\mathbf{d}$ (from the sensor to each 3D point) to derive $\cos(\alpha)=|\mathbf{d}\cdot\mathbf{n}|$, where $\mathbf{n}$ is the unit normal. These parameters are then applied to $I_{\text{m}}$ using Eq. (10), enabling the model to perform intensity compensation. To train the CNN module in the DCM, we use a dataset of 300 manually selected, well-calibrated compensated images as supervision data, optimizing the network with an $L_{1}$ loss function. We specifically selected samples with complete scene structures and clear textures while filtering out inputs containing excessive clutter (e.g., vegetation) or dynamic artifacts. The input to the CNN includes the original densified intensity image, as well as the per-pixel distance and incidence angle, allowing the CNN to dynamically predict per-pixel compensation parameters.
IV Experiment
We conducted a series of experiments to evaluate the effectiveness and robustness of the proposed method, focusing on the real-time LiDAR intensity image densification and its applications in loop closure detection and traffic lane detection. Data collection and real-world experiments were conducted using a Livox MID-360, mounted on a wheeled ground robot equipped with an NVIDIA RTX 3090 GPU and an Intel i7-1165G7 CPU. The experiments covered diverse scenarios, including varying lighting conditions (daylight, overcast, nighttime) and urban environments, to comprehensively evaluate the proposed method.
IV-A Evaluation of LiDAR Intensity Image Densification
@@
Since there are limited public datasets or prior works specifically addressing LiDAR intensity image densification, we trained several depth completion models, included traditional interpolation methods for comparison[26, 7, 18, 40], and reproduced LiDAR-Net [3], a method specifically designed for LiDAR intensity image completion, providing only quantitative results due to its non-open-source nature. An ablation study was conducted to analyze the contribution of the Adaptive Fusion Module (AFM). Table I presents the quantitative results, while Fig. 8 demonstrates the qualitative superiority of our method. Our method achieves high accuracy (highest PSNR/SSIM, lowest RMSE/MAE) while maintaining competitive runtime. Notably, as Marigold-dc[40] is a zero-shot diffusion-based method designed for depth completion, its performance is less effective in our task context; therefore, we only provide its quantitative results.
To evaluate the performance of our method in generating dense intensity images during robot motion, we tested on data (5-frame accumulation) captured when robot was subject to aggressive motion poses on uneven road. The resulting PSNR values (25.75–26.82) demonstrate the robustness of our approach against alignment noise and non-linear motion artifacts (Fig. 9). Additionally, we compared real-time results at various robot speeds in the supplementary material.
@@ IV-B1 Loop Closure Detection
To validate the effectiveness of our method, we adopt the approach of[34] but replace the high-resolution intensity data acquired by high-cost LiDAR with the dense intensity images generated by our method using Livox MID-360. Specifically, [34] extracts ORB feature descriptors from intensity images for loop closure detection. These descriptors are then represented as bag-of-words vectors for querying loop closures. After retrieving a candidate match, ORB descriptor matching is performed, and outliers are removed using PnP RANSAC[9] for verification. We conducted experiments in three distinct environments under two lighting conditions: daytime and nighttime (Fig. 10), using data collected under varying conditions to assess the generalization capability of our approach. Because the adopted loop closure detection method[34] is not based on deep learning, for a fair comparison, we compared our loop closure detection results with other non-deep-learning methods, including [45, 42], which are representative LiDAR-descriptor based methods. The results in Table II demonstrate that our method consistently outperforms other approaches, maintaining reliable performance despite changes in illumination.